# A Weight-aware-based Multi-source Unsupervised Domain Adaptation Method for Human Motion Intention Recognition

### v1

@@ IV Connecting Theory and Algorithm
Motivated by the upper bound of the expected error given in Theorem 3, we bridge theory and algorithm for MUDA and develop a new theoretical-ground MUDA algorithm, named WMDD, and then the detailed training steps of WMDD is givenare in this section.
@@ IV-A Theoretical Optimization Goal
Given hypothesis space $\mathcal{F}$ and samples $\mathcal{D}_{sj},\mathcal{D}_{t}$, margin $\mu$, the ideal error $\beta$ and complexity terms are assumed to be fixed. Therefore, minimizing the target error $\mathcal{L}_{Q}(h_{\bm{f}})$ in Theorem 3 is equivalent to the below optimization problem.
@@
then by substituting Eqs. (3) and (5) into Eq. (11), we have
$\min_{\bm{f}}\quad\sum_{j=1}^{N}\alpha_{j}\Big{\{}\mathbb{E}_{(\bm{x},y)\sim\widehat{P}_{j}}\big{(}\mathbb{W}_{\mu}\big{[}\omega_{\bm{f}}(\bm{x},y)\big{]}\big{)}$ (12)
$+$ $\max_{\bm{f^{\prime}}_{j}\in\mathcal{F}}\big{(}\mathbb{E}_{\widehat{Q}}\big{[}\mathbb{W}_{\mu}\big{(}\omega_{\bm{f^{\prime}}_{j}}\big{)}\big{]}-\mathbb{E}_{\widehat{P}_{j}}\big{[}\mathbb{W}_{\mu}\big{(}\omega_{\bm{f^{\prime}}_{j}}\big{)}\big{]}\big{)}\Big{\}}.$
The above optimization is a minmax game problem. This problem can be regarded as stackelberg game model, where auxiliary scoring function $\bm{f^{\prime}}_{j}$ is set as a leader and $\bm{f}$ is a follower. The form of optimization is given below.
@@
where the maximization of $\mathcal{P}_{\widehat{P}_{j},\widehat{Q}}^{\mu}=\mathbb{E}_{\widehat{Q}}\big{(}\mathbb{W}_{\mu}\big{)}-\mathbb{E}_{\widehat{P}_{j}}\big{(}\mathbb{W}_{\mu}\big{)}$ is the estimated MDD. In the Stackelberg game, the auxiliary function $\bm{f^{\prime}}_{j}$ is first updated, and then $\bm{f}$ is updated based on the result of $\bm{f^{\prime}}_{j}$. The key of the next step is to determine the value of $\bm{\alpha}$. Intuitively, the source domain with a small gap to the target domain should have a large weight. Therefore, $\bm{\alpha}$ can be defined as
@@ IV-B Bridging Theory and Algorithm
There is a gap between theory and algorithm since it is impractical to solve the above optimization problem directly. Firstly, considering the complexity of input space $\mathcal{X}$, we utilize the feature extractor $\phi(\cdot)$ to map the input space $\mathcal{X}$ into feature space $\Omega=\{\phi(\bm{x})~{}|~{}\bm{x}\in\mathcal{X}\}$. Then label function can be again defined as $h_{\bm{f}}:\phi(\bm{x})\rightarrow\arg\max_{y\in\mathcal{Y}}f_{y}(\phi(\bm{x}))$.
Secondly, since the ramp loss used in margin loss easily causes gradient vanishing, it is difficult to optimize Eq. (13) through stochastic gradient descent (SGD) [24]. Therefore, we want to find a surrogate function, denoted as $\mathcal{T}(\bm{f}(\phi(\bm{x})),y)$ here, for margin loss, which can be easily trained by SGD and keep the main property of the margin. Then, the two terms in Eq. (13) can be written as
@@
Therefore, the optimization problem (13) can be stated as two-stage optimization problem:
@@
$\min_{\phi,\bm{f}}\quad\mathcal{L}_{\bm{f},\phi}=\sum_{j=1}^{N}\alpha_{j}\Big{\{}\mathcal{L}_{{\widehat{P}}_{j}}^{\mu}\left(\bm{f}\right)+\mathcal{P}_{\widehat{P}_{j},\widehat{Q}}^{\mu}\left({\bm{f}},\bm{f^{\prime}}_{j}\right)\Big{\}},$ (19)
where $\mathcal{L}_{{\widehat{P}}_{j}}^{\mu}\left(\bm{f}\right)=\mathbb{E}_{(\bm{x},y)\sim\widehat{P}_{j}}\left\{-\log\big{[}\Theta_{y}\big{(}\bm{f}(\phi(\bm{x}))\big{)}\big{]}\right\}$. During the iterative optimization process, the auxiliary function $\bm{f^{\prime}}_{j}$ is firstly updated and then scoring function $\bm{f}$ and feature extractor $\phi$ are updated.
@@
The proof of Theorem 4 can be found in Appendix -D. Since the Kullback–Leibler divergence between two distributions is always non-negative and zero only when they are equal, $\gamma\log\gamma-(1+\gamma)\log(1+\gamma)$ is the global minimum of MDD and that the only solution is $P=Q$. This indicates the different choices of $\gamma$ do not result in the mismatch between source distribution $P$ and target distribution $Q$. When the $\gamma$ is fixed, the estimated MDD can effectively reflect the gap between $P$ and $Q$.
@@ IV-C Adversarial Learning Between Generator and Classifier
In practical implementation, feature extractor $\phi(\cdot)$ is approximated by a convolution neural network (CNN). The scoring function $\bm{f}$ and auxiliary function $\bm{f^{\prime}}_{j}$ $(j=1,...,N)$ are approximated by a fully connected neural network (NN). The $\bm{f}$ and $\bm{f^{\prime}}_{j}$ have the same network frame. In the following, we denote $\phi(\cdot)$ as the feature generator, $\bm{f}$ as the classifier, and $\bm{f^{\prime}}_{j}$ as the auxiliary classifier.
Due to the high real-time requirements of human intent recognition in practical applications, a lightweight NN is used here. Zhang et al.[17] pointed out that lightweight NN may have a better generalization ability, but may not fit the training dataset as accurately as a large NN. The ensemble method is commonly used to solve the above problem since it can enhance weak learners to make precise predictions [50].
@@
where $\bm{F}(\cdot)=\{\bm{f}^{(1)}(\cdot),...,\bm{f}^{(K)}(\cdot)\}$ and $\|\cdot\|_{1}$ is $L_{1}$ norm. Then, motivated by generative adversarial network (GAN) [51], adversarial learning (two-player zero-sum game) is introduced to further update feature generator $\phi$ and classifiers $\{\bm{f}^{(i)}\}_{i=1}^{K}$, described as
@@
In adversarial learning (two-player game), the ensemble classifiers are updated firstly fixing the feature generator, and then the feature generator is trained. The fitting ability and generalization ability can achieve a balance through constant learning.
@@ IV-D The Details of Algorithm
According to the above analysis, we propose a new Weighted multi-source UDA algorithm based on Margin Disparity Discrepancy (WMDD). To further improve the generalization of lightweight networks, the idea of adversarial learning is introduced to training steps. The detailed training steps are given below.
  • Step 1: The $j$-th auxiliary classifier $\bm{f^{\prime}}_{j}$ is optimized to maximize discrepancy between $j$-th source distribution $P_{j}$ and target distribution $Q$ defined in Eq. (18).
  • Step 2: The source domain weight $\bm{\alpha}$ is obtained by estimated MDD according to Eq. (14).
  • Step 3: The feature generator $\phi$ and classifiers $\{\bm{f}^{(i)}\}_{i=1}^{K}$ are optimized to minimize generalization error (target source) $\mathcal{L}_{\bm{f},\phi}$ in Eq. (19).
  • Step 4: Fixing feature generator $\mathcal{L}_{\phi}$, then classifiers $\{\bm{f}^{(i)}\}_{i=1}^{K}$ are updated by maximizing classifier discrepancy through Eq. (23).
  • Step 5: Fixing classifiers $\{\bm{f}^{(i)}\}_{i=1}^{K}$, then feature generator $\phi$ is updated by minimizing classifier discrepancy $\mathcal{L}_{\phi}$ according to Eq. (24).
Algorithm 1 gives the pseudo-code of the proposed method WMDD. The diagram of network updating can be found in Fig. 1. The final trained WMDD method can use a feature generator and ensemble classifiers to classify the target subject intention accurately.
@@ V-A Experimental Setup
For the ENABL3S dataset [52], ten subjects are invited to walk on several terrains and switch locomotion modes between sitting, standing, level ground walking, stair ascent, stair descent, ramp ascent, and ramp descent. It contains 7 class motion intentions in total, that is the above 7 activities. Each subject is asked to repeat walking on a circuit ten times. ENABL3S contains filtered EMG, IMU and joint angle signals, which are segmented by a 300 ms wide sliding window, and the segmented signals are used to recognize human motion intention [17].
@@
Experimental Details: In MUDA experiments, a target subject (domain) is selected first and then the remaining subjects are used as source subjects. ENABL3S and DSADS datasets contain 22,000 and 9000 signal segments, respectively. The data from each subject are randomly shuffled and divided into a training set (70%) and a test set (30%). Table I gives the base hyper-parameters of the WMDD method. The learning rate is set as $2\times 10^{-4}$, the optimizer is Adam, the number of iterations is $400$ and the batch size is set as $256$. The margin coefficient $\gamma$ is set to $0.1$. Note that WMDD sets the same parameters for ENABL3S and DSADS datasets, which is different from previous methods that need setting different parameters for different datasets. This is also the advantage of the proposed method WMDD.
The network framework information is as follows. The network is composed of one feature generator, $K$ classifiers and $N$ auxiliary classifiers, where the feature generator contains 3 convolution layers and the classifier includes 3 fully connected layers. The performance of WMDD has a close connection with domain weight $\bm{\alpha}$, number of ensemble classifiers $K$ and margin coefficient $\gamma$. The related experiments can be found in part C in this section. The code for WMDD is available at github.com/xiaoyinliu0714/WMDD. The code is run under an Intel(R) Xeon(R) Gold 6348 CPU @ 2.60GHz and an NVIDIA GeForce RTX 4090 GPU.
@@ V-B Feature Alignment and Classification Results
The Results of Feature Alignment: The dimensional reduction method t-SNE [54], which can keep the clustering of high dimensional space, is used to visualize the results of feature extraction following previous work [17]. Fig. 3 shows the t-SNE projection of the non-adapted input features and the hidden features adapted by WMDD. It can be found that the input features are not clustered, and the source distribution is not aligned with the target distribution. After feature alignment, the features of the same class almost cluster together for two datasets. For DSADS dataset, source and target features that belong to the same class are almost in the same region. However, for ENABL3S dataset, some source and target features that belong to the same class aren’t in the same region, which causes the classification accuracy of ENABL3S dataset is worse than that of DSADS dataset. The reason is that about 30% of data for ENABL3S are transition activities instead of steady activities, where the activity is defined as a transition activity if the locomotion mode of the current activity is different from that of the last activity.
TABLE II: The detailed results for different target subjects on ENABL3S and DSADS datasets. Std indicates the standard deviation. S1 represents the subject 1, which means the first subject is the target subject and other subjects are source subjects.
Dataset S1111 S2222 S3333 S4444 S5555 S6666 S7777 S8888 S9999 S10101010 Means Std
ENABL3S $97.8$ $95.3$ $94.9$ $91.5$ $94.6$ $94.2$ $91.3$ $95.4$ $96.7$ $94.5$ $\mathbf{94.6}$ $2.0$
DSADS $99.1$ $99.7$ $98.8$ $99.7$ $99.4$ $99.4$ $98.2$ $97.7$ $\mathbf{99.0}$ $0.7$
@@
TABLE IV: The ablation experiments for domain weight and adversarial learning. No-W and No-A represent that the domain weight method and adversarial learning are not employed, respectively. No-W-A represents both methods are not used.

### v2

@@ IV Connecting Theory and Algorithm
Based on the upper bound of the expected error given in Theorem 3, a novel theoretical-ground multi-source UDA algorithm is proposed in this section. We bridge theory and algorithm for MUDA, and elaborate on how to use lightweight networks to approximate the various terms of the above theory. Then, the detailed training steps of WMDD and the algorithm description are provided in this part.
@@ IV-A Theoretical Optimization Goal
The goal of optimization is to minimize the target error $\mathcal{L}_{Q}(h_{\bm{f}})$ in Theorem 3, that is minimizing the upper bound of generalization error. Given hypothesis space $\mathcal{F}$ and samples $\mathcal{D}_{sj}$ and $\mathcal{D}_{t}$ and margin $\mu$, the ideal error $\beta$ and complexity terms are assumed to be fixed. Therefore, the optimization goal can be written as:
@@
where $\mathcal{L}_{{\widehat{P}}_{j}}^{\mu}(\bm{f})$ denotes the expected error over the $j$-th source domains, and $d^{\mu}_{\bm{f}}(\widehat{P}_{j},\widehat{Q})$ is the margin disparity discrepancy between the $j$-th source and target distributions. Then by substituting Eqs. (3) and (5) into Eq. (10), we have
$\min_{\bm{f}}\quad\sum_{j=1}^{N}\alpha_{j}\Big{\{}\underbrace{\mathbb{E}_{(\bm{x},y)\sim\widehat{P}_{j}}\big{(}\mathbb{W}_{\mu}\big{[}\omega_{\bm{f}}(\bm{x},y)\big{]}\big{)}}_{{\color[rgb]{0,0,0}\text{Expected loss in $j$-th source: }\mathcal{L}_{{\widehat{P}}_{j}}^{\mu}\left(\bm{f}\right)}}$ (11)
$+$ $\underbrace{\max_{\bm{f^{\prime}}_{j}\in\mathcal{F}}\big{(}\mathbb{E}_{\widehat{Q}}\big{[}\mathbb{W}_{\mu}\big{(}\omega_{\bm{f^{\prime}}_{j}}\big{)}\big{]}-\mathbb{E}_{\widehat{P}_{j}}\big{[}\mathbb{W}_{\mu}\big{(}\omega_{\bm{f^{\prime}}_{j}}\big{)}\big{]}\big{)}}_{{\color[rgb]{0,0,0}\text{Margin disparity discrepancy: }d^{\mu}_{\bm{f}}\left(\widehat{P}_{j},\widehat{Q}\right)}}\Big{\}}.$
The above optimization problem is a minmax game problem. The form of optimization can be reformed as:
@@
where the maximization of $\mathcal{P}_{\widehat{P}_{j},\widehat{Q}}^{\mu}=\mathbb{E}_{\widehat{Q}}\big{(}\mathbb{W}_{\mu}\big{)}-\mathbb{E}_{\widehat{P}_{j}}\big{(}\mathbb{W}_{\mu}\big{)}$ is the estimated MDD. In this optimization problem, the auxiliary function $\bm{f^{\prime}}_{j}$ is first updated, and then $\bm{f}$ is updated based on the result of $\bm{f^{\prime}}_{j}$. The key of the next step is to determine the value of $\bm{\alpha}$. Intuitively, the source domain with a small gap to the target domain should have a large weight. Therefore, $\bm{\alpha}$ can be defined as
@@ IV-B Bridging Theory and Algorithm
There is a gap between theory and algorithm since it is impractical to solve the above optimization problem directly. Firstly, due to the complexity of input space $\mathcal{X}$, we utilize the feature extractor $\phi(\cdot)$ to map the high input space $\mathcal{X}$ into low feature space $\Omega=\{\phi(\bm{x})~{}|~{}\bm{x}\in\mathcal{X}\}$. Then label function can be again defined as $h_{\bm{f}}:\phi(\bm{x})\rightarrow\arg\max_{y\in\mathcal{Y}}f_{y}(\phi(\bm{x}))$.
Secondly, since the ramp loss used in the margin loss is prone to causing gradient vanishing, it is challenging to optimize Eq. (12) using stochastic gradient descent (SGD) [32]. Therefore, we aim to find a surrogate function, denoted as $\mathcal{T}(\bm{f}(\phi(\bm{x})),y)$, for the margin loss. This surrogate function should be easily trainable using SGD while retaining the key properties of the margin loss. Then, the two terms in Eq. (12) can be written as
@@
Therefore, the optimization problem in Eq. (12) can be written as the following two parts of optimization problem:
@@
$\min_{\phi,\bm{f}}\quad\mathcal{L}_{\phi,\bm{f}}=\sum_{j=1}^{N}\alpha_{j}\Big{\{}\mathcal{L}_{{\widehat{P}}_{j}}^{\mu}\left(\bm{f}\right)+\mathcal{P}_{\widehat{P}_{j},\widehat{Q}}^{\mu}\left({\bm{f}},\bm{f^{\prime}}_{j}\right)\Big{\}},\quad$ (18)
where $\mathcal{L}_{{\widehat{P}}_{j}}^{\mu}\left(\bm{f}\right)=\mathbb{E}_{(\bm{x},y)\sim\widehat{P}_{j}}\left\{-\log\big{[}\Theta_{y}\big{(}\bm{f}(\phi(\bm{x}))\big{)}\big{]}\right\}$. During the iterative optimization process, the auxiliary function $\bm{f^{\prime}}_{j}$ is firstly updated, and then scoring function $\bm{f}$ and feature extractor $\phi$ are updated.
@@
The proof of Theorem 4 can be found in Appendix VII-D. Since the Kullback–Leibler divergence between two distributions is always non-negative and zero only when they are equal, $\gamma\log\gamma-(1+\gamma)\log(1+\gamma)$ is the global minimum of MDD and that the only solution is $P=Q$. This indicates the different choices of $\gamma$ do not result in the mismatch between source distribution $P$ and target distribution $Q$. When the $\gamma$ is fixed, the estimated MDD can effectively reflect the gap between $P$ and $Q$.
@@ IV-C Adversarial Learning Between Generator and Classifier
In practical implementation, feature extractor $\phi(\cdot)$ is approximated by a convolution neural network (CNN). The scoring function $\bm{f}$ and auxiliary function $\bm{f^{\prime}}_{j}$ $(j=1,...,N)$ are approximated by a fully connected neural network (NN). The $\bm{f}$ and $\bm{f^{\prime}}_{j}$ have the same network frame. In the following part, we denote $\phi(\cdot)$ as the feature generator, $\bm{f}$ as the classifier, and $\bm{f^{\prime}}_{j}$ as the auxiliary classifier.
Due to the high real-time requirements of human intention recognition in practical applications, a lightweight neural network is used here. Zhang et al.[6] pointed out that lightweight NN may have a better generalization ability, but may not fit the training dataset as accurately as a large NN. The ensemble method is commonly used to solve the above problem since it can enhance weak learners to make precise predictions [40].
@@
where $\bm{F}(\cdot)=\{\bm{f}^{(1)}(\cdot),...,\bm{f}^{(K)}(\cdot)\}$ and $\|\cdot\|_{1}$ is $L_{1}$ norm. Then, motivated by the generative adversarial network (GAN) [5], adversarial learning is introduced to further update feature generator $\phi$ and classifiers $\{\bm{f}^{(i)}\}_{i=1}^{K}$ to improve the generalization ability, described as
@@
In adversarial learning, the ensemble classifiers are updated firstly fixing the feature generator, and then the feature generator is trained. The fitting ability and generalization ability can achieve a balance through constant learning.
Figure 3: The training steps of the proposed algorithm WMDD. $\bm{f}^{\prime}$, $\phi$ and $\bm{f}$ represent auxiliary classifiers, feature generator and classifier, respectively. Steps 1 to 4 are repeated in sequence until the training is completed.
@@ IV-D The Details of Algorithm
According to the above analysis, we propose a new Weighted multi-source UDA algorithm based on Margin Disparity Discrepancy (WMDD). Fig. 3 shows the detailed training steps of algorithm WMDD. It contains 4 steps. First, $N$ auxiliary classifiers are trained to estimate MDD, and then the source domain weights $\alpha$ are calculated. Second, the feature generator and $K$ classifiers are optimized based on minimizing generalization error. The last two steps are adversary learning, and the feature generator and $K$ classifiers are updated separately. Steps 1 to 4 are repeated in sequence until the training is completed. The detailed illustrations are given below.
  • Step 1: The $j$-th auxiliary classifier $\bm{f^{\prime}}_{j}$ is optimized to maximize the discrepancy defined in Eq. (17). Then the source domain weight $\bm{\alpha}$ is obtained by the estimated MDD according to Eq. (13).
  • Step 2: The feature generator $\phi$ and classifiers $\{\bm{f}^{(i)}\}_{i=1}^{K}$ are optimized to minimize generalization error (target source) $\mathcal{L}_{\bm{f},\phi}$ in Eq. (18).
  • Step 3: Fixing feature generator $\phi$, then classifiers $\{\bm{f}^{(i)}\}_{i=1}^{K}$ are updated by maximizing classifier discrepancy through Eq. (22).
  • Step 4: Fixing classifiers $\{\bm{f}^{(i)}\}_{i=1}^{K}$, then feature generator $\phi$ is updated by minimizing classifier discrepancy $\mathcal{L}_{\phi}$ according to Eq. (23).
The state of the data flow is as follows: Initially, multimodal data undergoes feature extraction through a feature generator network. Subsequently, the extracted feature information diverges into two pathways. The first pathway processes the data through an auxiliary classifier network to determine the source domain weights, while the second pathway feeds the data into a classifier network to obtain prediction outcomes. The visualization of specific data flow can be found in Fig. 1. The training procedures for each component of the network are detailed in Algorithm 1. The final WMDD method can use a feature generator and ensemble classifiers to classify the target subject intention accurately.
@@ V-A Experimental Setup
For the ENABL3S dataset [9], ten subjects are invited to walk on several terrains and switch locomotion modes between sitting, standing, level ground walking, stair ascent, stair descent, ramp ascent, and ramp descent. It contains 7 class motion intentions in total, that is the above 7 activities. Each subject is asked to repeat walking on a circuit ten times. ENABL3S contains filtered EMG, IMU, and joint angle signals, which are segmented by a 300 ms wide sliding window, and the segmented signals are used to recognize human motion intention [6].
@@
Experimental Details: In MUDA experiments, a target subject (domain) is selected first and then the remaining subjects are used as source subjects. ENABL3S and DSADS datasets contain 22,000 and 9000 signal segments, respectively. The data from each subject are randomly shuffled and divided into a training set (70%) and a test set (30%). Table I gives the base hyper-parameters of the WMDD method. The learning rate is set as $2\times 10^{-4}$, the optimizer is Adam, the number of iterations is $400$ and the batch size is set as $256$. The margin coefficient $\gamma$ is set to $0.1$. The number of auxiliary classifiers $N$ is equal to the number of source subjects and set as $9$ and $7$ for ENABL3S and DSADS datasets, respectively. Note that WMDD sets the same parameters for ENABL3S and DSADS datasets, which is different from previous methods that need setting different parameters for different datasets. This is also the advantage of the proposed method WMDD.
The network framework information is as follows. The lightweight network is composed of one feature generator, $K$ classifiers, and $N$ auxiliary classifiers, where the feature generator contains 3 convolution layers and the classifier includes 3 fully connected layers. The number of network parameters is $2.21$ M and $1.62$ M for ENABL3S and DSADS datasets, respectively. The performance of WMDD has a close connection with domain weight $\bm{\alpha}$, number of ensemble classifiers $K$, and margin coefficient $\gamma$. The related experiments can be found in part C in this section. The code for WMDD is available at github.com/xiaoyinliu0714/WMDD. The code is run under an Intel(R) Xeon(R) Gold 6348 CPU @ 2.60GHz and an NVIDIA GeForce RTX 4090 GPU.
@@ V-B Feature Alignment and Classification Results
The Results of Feature Alignment: The dimensional reduction method t-SNE [15], which can keep the clustering of high dimensional space, is used to visualize the results of feature extraction following previous work [6]. Fig. 4 shows the t-SNE projection of the non-adapted input features and the hidden features adapted by WMDD. It can be found that the input features are not clustered, and the source distribution is not aligned with the target distribution. After feature alignment, the features of the same class almost cluster together for two datasets. For the DSADS dataset, source and target features that belong to the same class are almost in the same region. However, for the ENABL3S dataset, some source and target features that belong to the same class aren’t in the same region, which causes the classification accuracy of the ENABL3S dataset to be worse than that of the DSADS dataset. The reason is that about 30% of data for ENABL3S are transition activities instead of steady activities, where the activity is defined as a transition activity if the locomotion mode of the current activity is different from that of the last activity.
TABLE II: The detailed results for different target subjects and corresponding training and testing times for ENABL3S and DSADS datasets. Std indicates the standard deviation. S1 represents the subject 1, which means the first subject is the target subject and other subjects are source subjects.
Dataset S1111 S2222 S3333 S4444 S5555 S6666 S7777 S8888 S9999 S10101010 Means Std Training Testing
[…]
@@
The Classification Results on ENABL3S and DSADS: To answer $\mathbf{Q_{1}}$, we compare different domain adaptation methods in ENABL3S and DSADS datasets, including DANN [13], MMD [12], MCD [11], DFA [14], GFA [7] and EDHKD [6], where EDHKD is the state-of-the-art method for HMI using domain adaptation method. The classification results on ENABL3S and DSADS of these methods come from article [6]. Table II gives the detailed results for different target subjects on ENABL3S and DSADS datasets, and shows the related training and testing times of lightweight network. It indicates that the performance of WMDD varies slightly across different target subjects, but it can achieve a relatively high average performance. Additionally, the training process of the WMDD requires $15$ and $11$ minutes on ENABL3S and DSADS datasets, indicating low computational cost. Moreover, the testing time of for a single sample is about $0.5$ ms, which can guarantee the real-time of the testing.
Table III compares the mean accuracy of classification on the target subject for different domain adaptation methods. It shows that WMDD achieves the accuracy of $94.6\%\pm 2.0\%$ and $99.0\%\pm 0.7\%$ for target subject in ENABL3S and DSADS datasets respectively, which are $0.2\%$ and $1.6\%$ higher than the state-of-the-art HMI result using domain adaptation methods. This verifies that WMDD can achieve a better performance than previous methods in both ENABL3S and DSADS datasets. However, the improvement in ENABL3S is not significant despite that the samples of ENABL3S are larger than DSADS. The reason lies in the effects of transition activities. The results of classification are consistent with that of feature alignment. In addition, we give the final source domain weight $\bm{\alpha}$ when the first subject is the target subject. The weights are $0.1108$, $0.1106$, $0.1108$, $0.1106$, $0.1110$, $0.1118$, $0.1121$, $0.1114$, $0.1108$ for ENABL3S dataset and $0.1438$, $0.1431$, $0.1417$, $0.1454$, $0.1441$, $0.1396$, $0.1424$ for DSADS dataset.
The superior performance of our method compared to EDHKD can be attributed to two key factors: 1) EDHKD does not adequately account for the variations among individual source subjects, which may negatively impact classification accuracy. In contrast, our approach explicitly considers these inter-source differences, leading to improved classification outcomes. 2) We extend the margin disparity discrepancy (MDD) to the context of multi-source unsupervised domain adaptation. MDD incorporates more comprehensive generalization bound information, where the generalization bound is affected by the margin parameter $\mu$. By selecting an appropriate $\mu$, the classification error on the target subject can be minimized, thereby enhancing performance on the target domain.
V-C Ablation Experiments
Effects of Weight $\bm{\alpha}$ and Adversarial Learning: To answer $\mathbf{Q_{2}}$, the ablation experiments for two parts are designed (see Table V). The weight $\bm{\alpha}$ reflects the discrepancy between different source domains and target domain. The adversarial learning is used to improve the classifier’s performance. From this table, the average performance in ENABL3S and DSADS is improved by $1.07\%$ and $1.02\%$ respectively compared with method No-W. The average performance is improved by $1.72\%$ and $7.25\%$ respectively compared with method No-A, and the maximum performance improvement reaches $5.91\%$ on one of the target subject. Therefore, the usage of domain weight and adversarial learning can both improve the performance of the classifier, and the effect of adversarial learning is more significant.
TABLE IV: The relationship between the classification accuracy and the number of classifiers $K$, and the relationship between classification accuracy and trade-off coefficient $\eta$ on ENABL3S and DSADS datasets.000
Number of classifiers $K$ $1$ $3$ $5$ $7$ $9$ $11$ $13$ $15$ $17$ $19$ $21$ $23$ $25$
ENABL3S $93.3$ $91.8$ $\mathbf{97.8}$ $96.4$ $96.5$ $96.5$ $95.4$ $95.9$ $96.2$ $95.6$ $95.8$ $96.1$ $96.4$
DSADS $96.8$ $98.0$ $\mathbf{99.1}$ $98.2$ $97.7$ $98.0$ $98.5$ $98.0$ $\mathbf{99.1}$ $97.4$ $98.2$ $98.2$ $97.4$
Trade-off coefficient $\eta$ $1$ $2$ $3$ $4$ $5$ $7$ $8$ $9$ $11$ $12$ $13$ $14$ $15$
ENABL3S $96.4$ $93.6$ $96.7$ $\mathbf{97.6}$ $\mathbf{97.8}$ $97.0$ $96.5$ $97.3$ $96.1$ $96.4$ $96.4$ $94.7$ $94.0$
DSADS $94.4$ $98.5$ $97.4$ $99.1$ $\mathbf{99.4}$ $99.1$ $99.1$ $99.1$ $97.7$ $98.5$ $96.2$ $98.2$ $92.1$
TABLE V: The ablation experiments for domain weight and adversarial learning. No-W and No-A represent that the domain weight method and adversarial learning are not employed, respectively. No-W-A represents both methods are not used.